OpenStudy (mikezack123):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is.

OpenStudy (radar):

There are numerous ways to solve this, this is just one of many. C = 5/9 (F-32) Multiply the equation by 9/5 getting 9/5 C = F-32 Add 32 to both sides of equal sign getting 9/5 C + 32 = F There you are F=9/5 C + 32

OpenStudy (radar):

Just for fun see what the temperature in fahrenheit (F) is when it is -40 degrees Centigrade.

OpenStudy (radar):

F =9/5 (-40) + 32 F = -72 +32 F= -40 The one temperature where they read the same.

OpenStudy (mikezack123):

really? -40 F is the same as -40C?

OpenStudy (radar):

Yes, the equations show that, and math never lies.

OpenStudy (mikezack123):

LOL, wow, thats a great fact, i never knew that. thanks! :)

OpenStudy (radar):

Good luck with your studies, and now you can convert from one system to another.

OpenStudy (mikezack123):

:)